write the function form of 4x-2y=18
Answer by funmath(2933) 
 write the function form of 4x-2y=18
-4x+4x-2y=-4x+18
0-2y=-4x+18
-2y=-4x+18
-2y/-2=-4x/-2+18/-2
y=2x-9 <--in some books this is enough, in others, they replace y with f(x)
f(x)=2x-9
I hope one of these is what your book wants.
Happy Calculating!!!
|
|
|
| |